4. What would be the shape of:



(a) The IS curve if investment does not depend on the interest rate?



(b) The LM curve if the demand for money does not depend on interest rate?



(c) The LM curve if the demand for money does not depend on income?

1
Expert's answer
2021-12-29T13:21:29-0500

a)     If investment does not depend on the interest rate, the IS curve is vertical. The IS curve depicts the relationship that exists between the interest rate and the level of income in a market for goods and services that is in equilibrium. It describes the income and interest rate combinations that meet the equation Y = C(Y-T) + I(r) + G. Nothing in the IS equation depends on the interest rate if investment does not; income must adapt to ensure that the amount of goods produced, Y, equals the quantity of goods demanded, C + I + G. As a result, at this level, the IS curve is vertical. Because the IS curve determines Y, monetary policy has no effect on production. Only the interest rate can be influenced by monetary policy. Fiscal policy, on the other hand, is effective: output rises by the same amount that the IS curve shifts.

b)     If money demand does not depend on the interest rate, the LM curve is vertical. The LM curve depicts the income and interest rate combinations at which the money market is in equilibrium. If money demand is unaffected by interest rates, the LM equation can be written as M/P = L. (Y). There is only one level of income at which the money market is in equilibrium for any given level of real balances M/P. The LM curve is thus vertical. Fiscal policy can no longer affect output; it can only affect the interest rate. Monetary policy works: a movement in the LM curve boosts output by the same amount as the shift.

c)     If money demand does not depend on income, the LM curve is horizontal. If money demand is independent of income, the LM equation can be written as M/P = L. (r). There is only one interest rate level at which the money market is in equilibrium for every given level of real balances M/P. The LM curve is thus horizontal. Fiscal policy is particularly effective: output rises in proportion to the shift in the IS curve. Monetary policy is also effective: expanding the money supply lowers the interest rate, causing the LM curve to shift downward.